OpenStudy (anonymous):

what is the value of x when 10(x+2) =5(x+8) ? best response gets a medal !!!!

OpenStudy (anonymous):

10x +20 = 5x +40 -5x -5x 5x +20 = 40 -20 -20 5x/5 = 20/5 x = 4

OpenStudy (anonymous):

thank u !!

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

What you want to do, is find a number that you can plug in for X that will make both of the equations equal; meaning both equations will have the same answer. Here's some examples, Lets plug in 2 for X: 10(2+2)=5(2+8) You start from left to right and begin in the parenthesis: 2+2=4 10(4)=40 2+8=10 10(10)=100 4 does not equal 100 Let's try 4 for X: 4+2=6 10(6)=60 4+8=12 5(12)=60 60 does equal 60 4=x